# AAPM: Large Language Model Agent-based Asset Pricing Models

### v1

@@ Abstract
In this study, we propose a novel asset pricing approach, LLM Agent-based Asset Pricing Models (AAPM), which fuses qualitative discretionary investment analysis from LLM agents and quantitative manual financial economic factors to predict excess asset returns. The experimental results show that our approach outperforms machine learning-based asset pricing baselines in portfolio optimization and asset pricing errors. Specifically, the Sharpe ratio and average $|\alpha|$ for anomaly portfolios improved significantly by 9.6% and 10.8% respectively. In addition, we conducted extensive ablation studies on our model and analysis of the data to reveal further insights into the proposed method.
AAPM: Large Language Model Agent-based Asset Pricing Models
Junyan Cheng Thayer School of Engineering Dartmouth College Hanover, NH 03755, USA jc.th@dartmouth.edu          Peter Chin Thayer School of Engineering Dartmouth College Hanover, NH 03755, USA pc@dartmouth.edu
@@ 1 Introduction
The pricing of financial assets, such as stocks, has been a focal point in empirical financial economics research. It has a significant impact on social good by moving towards Pareto efficiency in capital allocation. Current asset pricing methods rely on carefully crafting manual macroeconomic indicators or company-specific factors as predictors of future excess returns Fama and French (1992, 2015). Despite its great success in the real-world market, they have been challenged by the Efficient Market Hypothesis (EMH) that manual factors will ultimately lose their predictive power in an efficient market when these predictors are fully discovered and used by market participants.
Due to this rationale, linguistic data, which are the primary sources of traditional discretionary investing, become essential. This is because the dynamics of society and the market are largely driven by the information flow of language. This is also evident in the real financial world, where discretionary portfolio management remains significant today Abis (2020). Such investment decisions are mainly shaped by the manager’s experience and intuition, as they evaluate assets and determine their value based on information from news, investigations, reports, etc., instead of depending on quantitative models.
@@
This phenomenon highlights two key points. First, qualitative discretionary analysis can uncover valuable pricing insights that are absent in economic indicators or market data. Second, even with the integration of current NLP and semantic analysis methods, quantitative factor models have not fully captured these insights. Achieving the synergy between both remains a complex yet appealing objective Cao et al. (2021). Nonetheless, leveraging linguistic information is complicated as it requires financial reasoning and long-term memory of tracking events and company impressions to interpret. Furthermore, suboptimal interactions in model design between linguistic and manual factors can end up as noise Bybee et al. (2023).
@@
Figure 1: The LLM agent produces analysis report from the latest news through a multi-step refinement, incorporating past reports and domain knowledge from memory. For simplicity, the filter for irrelevant news is excluded. A macro and micro note, continuously updated by the latest analysis report, is used to provide additional context. The average embedding of daily analysis reports will be input into the pricing network along with daily manual factors.
@@
In this study, we introduce a novel asset pricing approach, LLM Agent-based Asset Pricing Models (AAPM), which fuses discretionary investment analysis simulated by an LLM agent and quantitative factor-based methods. AAPM employs the LLM agent to iteratively analyze the latest news, supported by a memory of previous analysis reports and a knowledge base comprising books, encyclopedias, and journals. The embedding of analysis reports is merged with manual factors to predict future excess asset returns. Besides offering a performance edge, our method also provides enhanced interpretability through generated analysis reports. We evaluate our approach with a dataset consisting of two years of news and approximately 70 years of economic and market data. The experimental results show that our approach surpasses machine learning-based asset pricing baselines, achieving a 9.6% increase in the Sharpe ratio and a 10.8% improvement in the average $|\alpha|$ for asset pricing errors in character-section portfolios. Our primary contributions are summarized as follows:
@@
  • Proposed a hybrid asset pricing framework combines qualitative discretionary analysis and quantitative manual factors.
@@ 2 Related Work
2.1 Asset Pricing for Security
@@
Asset pricing aims to search for the fair price of financial assets, such as securities. Sharpe (1964) introduced the groundbreaking Capital Asset Pricing Model (CAPM), which breaks down the expected return of an asset into a linear function of the market return. Various extensions of the CAPM have been developed. Merton (1973) incorporated wealth as a state variable, while Lucas Jr (1978) considered consumption risk as a pricing factor. The single-factor CAPM was later expanded into multi-factor models. Fama and French (1992) proposed the Fama-French 3-factor (FF3) model, which explains returns by size, leverage, book-to-market equity, and earnings-price ratios. They later revised it to a 5-factor model Fama and French (2015). Furthermore, Carhart (1997) identified momentum as an additional factor. Ross (1976) formulated the Arbitrage Pricing Theory (APT), which considers asset pricing as an equilibrium in the absence of arbitrage opportunities. The Stochastic Discount Factor (SDF) calculates the price by discounting future cash flows using a stochastic pricing kernel Cochrane (2009).
2.2 Financial Machine Learning
@@
The application of machine learning techniques has been introduced to explore the non-linear interactions among the growing “factor zoo” Feng et al. (2020). Instrumented Principal Component Analysis (IPCA) was developed by Kelly et al. (2020) to estimate latent factors and their loadings from data. Gu et al. (2020) introduced a deep neural network to model interactions. Gu et al. (2021) proposed a conditional autoencoder that considers latent factors and asset characteristics as covariates. Chen et al. (2024) utilized Generative Adversarial Networks to train a neural SDF based on the methods of moments. Additionally, Bybee et al. (2021) conducted an analysis of the Wall Street Journals (WSJ) to gauge the state of the economy. Based on this analysis, Bybee et al. (2023) further suggested using Latent Dirichlet Allocation (LDA) to analyze monthly news topics from WSJ as pricing factors. Recent NLP methods Xu and Cohen (2018); Xie et al. (2022) have been employed to forecast stock movements, in contrast to asset pricing, they do not aim to find interpretable factors that explain anomalies in excess asset returns. Our LLM-based approach offers an alternative interpretation through analysis reports.
2.3 Large Language Model Agents
@@
LLM agents possess powerful emergent capabilities, such as reasoning, planning, and tool-using Achiam et al. (2023). The core of LLM agent programming lies in prompting, which employs contextual hinting text to regulate the output of LLM Liu et al. (2023). Several prompting strategies have been proposed. Chain-of-Thoughts (CoT) Wei et al. (2022) encourages the agent to reason in a step-by-step manner. Yao et al. (2022) introduced the ReAct prompt, enabling the agent to refine its output based on the results of previous attempts. It allows the agent to use external tools, such as databases and search engines. Memory is another crucial component of LLM agents. Hu et al. (2023) introduced databases as symbolic memories. Packer et al. (2023) stores dialogues in both long- and short-term memory, analogously to operating systems. Cheng and Chin (2024) developed an agent capable of making “investment” decisions on social science time series based on input news, reports, etc., and knowledge base, as well as the Internet. We focus on using the agent to simulate discretionary investment decision-making to synergize qualitative and quantitative asset pricing.
@@ 3 Method
Given a state vector $V_{\tau,a}$ at a time point $\tau\in\{0,1,2,$...\}, which represents the current status of the market, society, and an asset $a$, an asset pricing model predicts the excess returns $r_{\tau+1,a}$ of the asset at the subsequent time point, expressed as $P(r_{\tau+1,a}|V_{\tau,a})$. In our study, each time point corresponds to one day. In traditional factor-based methods, the state $V_{\tau,a}\in\mathcal{N_{F}}^{N}$ is a vector composed of $N_{F}$ factors that are manually derived from economic indicators, market data, asset characteristics, etc. For instance, the market excess return, the performance disparity between small and large firms, and the difference between high and low book-to-market companies in the Fama-French 3-factor model Fama and French (1992). Recently, Bybee et al. (2021) demonstrated that a collection of business news can serve as an alternative representation of macroeconomic conditions, while Bybee et al. (2023) employs LDA to extract news characteristics as economic predictors for pricing. Building on this idea, we use the average embedding of analysis reports that mine values from the news as a proxy for the society, economic, and market states.
Figure 2: Visualization of the key words in the titles of news articles on the days when the market return is positive (left) and when it is negative (right).
Business news in major media outlets like the WSJ carries important market information, however they typically restrict their interpretations and opinions, leaving room for discretionary analysis. It is crucial to understand that business events are often interrelated.
@@
Based on these observations, we introduce AAPM, utilizing an LLM agent with long-term memory of domain knowledge and historical news analysis to iteratively analyze the input news and generate the analysis report, as detailed in Section 3.1. Subsequently, we combine these qualitative analysis reports and quantitative manual factors to feed into our hybrid asset pricing network in Section 3.2.
Figure 3: The demonstration of our hybrid asset pricing network. The purple boxes mark the computational components. Yellow boxes are data, the circled plus symbol means contatenation. The MSE loss computed with predicted returns feedback to update the network.
@@ 3.1 Discretionary Analysis with LLM Agent
The agent utilizes the latest news $x_{t}$ at time $t$ (e.g., a WSJ article published at 9:32 AM on 6 June 2020), along with a note $n_{t}$ on macroeconomics and market trends, to generate an analysis report $R_{t}$. The note $n_{t}$ is initalized with a macroeconomics summary $n_{0}$ produced by GPT-3.5-Turbo-1106 Brown et al. (2020), the LLM used in our study, prior to its knowledge cut-off date $d_{k}$. It offers necessary macroscopic context on economic and societal trends not directly available from the news or the memory. The note is then iteratively updated to $n_{t^{\prime}}$ with the new analysis report $R_{t}$ to keep the context up-to-date, and we also prompt the agent to document investment ideas and market thoughts in the notes to provide a short-term background such as the trends on the market, long-term research oppurtunities to watch. To ensure the note is continuously updated without missing information while preventing information leakage, the dataset in our study starts from $d_{k}+1$, immediately following the knowledge cut-off date.
@@
The analysis process begins with generating a refined news item $x^{\prime}_{t}$ that summarizes key information from the raw input $x_{t}$. This step helps control the input length and standardizes the format and style. The refined news $x^{\prime}_{t}$ and the note $n_{t}$ are then combined to form an input $I_{t}$ for the agent. The agent will determine if the news contains investment information: if not, it will be skipped; otherwise, an initial analysis report $R_{t}^{0}$ will be created. The report undergoes iterative refinement over $N$ rounds. In each round $i$, the report $R_{t}^{i-1}$ is used to query an external memory $M^{t}$, a vector database initialized with the SocioDojo knowledge base Cheng and Chin (2024), which includes textbooks, encyclopedias, and academic journals in fields such as economics, finance, business, politics, and sociology. We use BGE Xiao et al. (2023) as the embedding model $f_{e}$, which maps text to a vector $e\in\mathcal{R}^{d_{emb}}$ for querying the memory. This choice is based on the MTEB leaderboard Muennighoff et al. (2022), where we selected the best retrieval model considering performance, model size, and embedding vector length. In each round $i$, the top-$K$ most relevant items $\{m^{t,i}_{j}\}_{j=1}^{K}\subset M^{t}$ are retrieved and provided to the agent along with the report $R^{i-1}_{t}$ to produce the refined report $R_{t}^{i}$. The report $R_{t}^{N}$ generated after the $N$-th round is used as the final analysis report $R_{t}$ for the news $x_{t}$ and to update the note as $n_{t^{\prime}}$. Then it is inserted into the memory $M^{t}$ for future reference and pricing, updating the memory to $M^{t^{\prime}}$.
@@
The pricing network will utilize the analysis reports $\{R_{t^{d}_{i}}\}_{i=1}^{N_{d}}$ of all filtered news $\{x_{t^{d}_{i}}\}_{i=1}^{N_{d}}$ for a given day $d$, where $N_{d}$ represents the number of filtered news items on day $d$. Figure 1 provides an overview of the entire analysis process. The prompts employed in our agent are detailed in Appendix E. In Section 4.4, we conduct experiments on our agent design and the impact of $N$ and $K$.
@@ 3.2 Hybrid Asset Pricing Network
where $\kappa(L,i)$ is an exponential decay kernel defined as $\frac{\eta^{L-i}}{\sum_{j=1}^{L}\eta^{L-j}}$. The decay coefficient is denoted as $0<\eta<1$. We form a raw hybrid state $h_{d,a}=[s_{d};v_{d,a}]$ by concatenating the smoothed daily state $s_{d}$ with a vector $v_{d,a}\in\mathcal{R}^{N_{F}}$ of $N_{F}$ manual-constructed financial economic factors. The asset $a$ is indexed by a permanent number (permno) from the Center for Research in Security Prices (CRSP)  [^1] https://www.crsp.org/ database. The hybrid state is subsequently downsampled by $h^{\prime}_{d,a}=\sigma(W_{S}h_{d,a})$, where $\sigma$ denotes the ReLU function and $W_{S}\in\mathcal{R}^{d_{model}\times(d_{emb}+N_{F})}$ is a parameter matrix.
To capture the asset-specific loading to the hybrid state especially to the asset-agnostic $s_{d}$, we define an asset embedding $E\in\mathcal{R}^{N_{A}\times d_{model}}$, which can be looked up via the permnos of the assets. Here, $N_{A}$ denotes the total number of assets and $d_{model}$ is the dimension of the embedding. We then concatenate the asset embedding $E_{a}$ with the downsampled hybrid state to form $\hat{h}_{d,a}=[h_{d,a}^{\prime};\sigma(E_{a})]$, the asset-specfic hybrid state for $a$.
@@
$\begin{split}\arg\min_{\theta}\frac{1}{N_{A}N_{D}}\sum_{d,a}(r_{d+1,a}-\hat{r}_{d+1,a})^{2},\\ where\ r_{d+1,a}=f_{H}(h_{d,a};\theta)\end{split}$
Where $N_{D}$ denotes the number of days in the training set. The model is trained for $T$ episodes with a batch size of $B$. We initially pre-train this hybrid asset pricing network $f_{H}$ to make use of the historical factor data available before the beginning of the news dataset. During this pre-training phase, a placeholder embedding (such as the embedding for the word "Null") is utilized.

### v2

# Empirical Asset Pricing with Large Language Model Agents

@@ Abstract
In this study, we introduce a novel asset pricing model leveraging the Large Language Model (LLM) agents, which integrates qualitative discretionary investment evaluations from LLM agents with quantitative financial economic factors manually curated, aiming to explain the excess asset returns. The experimental results demonstrate that our methodology surpasses traditional machine learning-based baselines in both portfolio optimization and asset pricing errors. Notably, the Sharpe ratio for portfolio optimization and the mean magnitude of $|\alpha|$ for anomaly portfolios experienced substantial enhancements of 10.6% and 10.0% respectively. Moreover, we performed comprehensive ablation studies on our model and conducted a thorough analysis of the method to extract further insights into the proposed approach. Our results show effective evidence of the feasibility of applying LLMs in empirical asset pricing.
@@ 1 Introduction
The pricing of financial instruments, including equities, has occupied a central position in empirical financial economics studies. This research exerts a notable influence on societal welfare by promoting Pareto efficiency in the distribution of capital. Present asset valuation techniques typically involve the meticulous development of manual macroeconomic indicators or factors specific to individual companies to forecast future excess returns Fama & French (1992; 2015). Although these methods have achieved considerable success in practical markets, they face criticism from the Efficient Market Hypothesis (EMH). The EMH suggests that in an efficient market, the predictive power of manually designed factors diminishes as market participants eventually uncover and utilize these predictors.
Due to this rationale, linguistic data, which are the primary sources of traditional discretionary investing, become essential. This is attributed to the substantial influence of linguistic information flow on societal and market dynamics. The importance of this is further underlined in actual financial settings, where discretionary portfolio management continues to play a crucial role Abis (2020). Investment decisions in this realm are predominantly informed by the manager’s expertise and intuition, with assets being assessed and appraised on the basis of data derived from news, inquiries, and reports, among other sources, rather than relying on quantitative modeling.
@@
This phenomenon emphasizes two significant aspects. First, qualitative discretionary analyses have the potential to reveal crucial pricing information that is not present in traditional economic indicators or market datasets. Second, despite the incorporation of current natural language processing and semantic analysis approaches, quantitative factor models have yet to fully integrate these insights. Realizing the synergy between these approaches remains a challenging yet enticing goal Cao et al. (2021). Nevertheless, the use of linguistic information is compounded by the need for financial reasoning and the ability to maintain a long-term memory of tracking events and company perceptions for accurate interpretation. Moreover, suboptimal interactions in model architecture between linguistic and manual elements may result in noise instead of information gain Bybee et al. (2023).
@@
Figure 1: The LLM agent produces analysis reports from the latest news through a multi-step refinement, incorporating past reports and domain knowledge from memory. For simplicity, the filter for irrelevant news is excluded. A macro and micro note, continuously updated by the latest analysis report, is used to provide additional context. The average embedding of daily analysis reports will be input into the pricing network along with daily manual factors.
Recent advancements in Large Language Models (LLMs) promote the utilization of textual and other alternative data in financial domains Cheng & Chin (2024). However, most existing methods focus on the trading strategies for individual stocks Zhang et al. (2024); Fatemi & Hu (2024), other than the macroscopic asset pricing problem which aims to discover the explanatory factors across all assets in the market rather than optimizing for the prediction on a single specific company.
@@
In this study, we introduce a novel LLM agent asset pricing model, which fuses discretionary investment analysis simulated by an LLM agent and quantitative factor-based methods. Our method employs the LLM agent to iteratively analyze the latest news, supported by a memory of previous analysis reports and a knowledge base comprising books, encyclopedias, and journals. The embedding of analysis reports is merged with manual factors to predict future excess asset returns. Besides offering a performance edge, our method also provides enhanced interpretability through generated analysis reports. We evaluate our approach with a dataset consisting of three years of news and approximately 70 years of economic and market data. The experimental results show that our approach surpasses machine learning-based asset pricing baselines, achieving a 10.6% increase in the Sharpe ratio for portfolio optimization and a 10.0% improvement in the average $|\alpha|$ for asset pricing errors in character-section portfolios. Our work demonstrates the viability of applying LLMs for asset pricing. The primary contributions are summarized as follows:
@@
  • Proposed a hybrid asset pricing framework that combines qualitative discretionary analysis and quantitative manual factors.
@@ 2 Related Work
Asset Pricing for Security.
@@
Asset pricing aims to search for the fair price of financial assets, such as securities, by discovering underlying explanatory factors across different assets. Sharpe (1964) introduced the groundbreaking Capital Asset Pricing Model (CAPM), which breaks down the expected return of an asset into a linear function of the market return. Various extensions of the CAPM have been developed. Merton (1973) incorporated wealth as a state variable, while Lucas Jr (1978) considered consumption risk as a pricing factor. The single-factor CAPM was later expanded into multi-factor models. Fama & French (1992) proposed the Fama-French 3-factor (FF3) model, which explains returns by size, leverage, book-to-market equity, and earnings-price ratios. They later revised it to a 5-factor model Fama & French (2015). Furthermore, Carhart (1997) identified momentum as an additional factor. Ross (1976) formulated the Arbitrage Pricing Theory (APT), which considers asset pricing as an equilibrium in the absence of arbitrage opportunities. The Stochastic Discount Factor (SDF) calculates the price by discounting future cash flows using a stochastic pricing kernel Cochrane (2009). We introduce LLMs to produce news-based macroeconomic factors for additional explanations of prices.
Financial Machine Learning.
@@
The application of machine learning techniques has been introduced to explore the non-linear interactions among the growing “factor zoo” Feng et al. (2020). Instrumented Principal Component Analysis (IPCA) was developed by Kelly et al. (2020) to estimate latent factors and their loadings from data. Gu et al. (2020) introduced a deep neural network to model interactions. Gu et al. (2021) proposed a conditional autoencoder that considers latent factors and asset characteristics as covariates. Chen et al. (2024) utilized Generative Adversarial Networks to train a neural SDF based on the methods of moments. Kelly et al. (2025) introduced a linear transformer which further reduces the pricing errors. Furthermore, Bybee et al. (2021) conducted an analysis of the Wall Street Journals (WSJ) to gauge the state of the economy. Based on this analysis, Bybee et al. (2023) further suggested using Latent Dirichlet Allocation (LDA) to analyze monthly news topics from WSJ as pricing factors. Recent NLP methods Xu & Cohen (2018); Xie et al. (2022) have been employed to forecast stock movements, in contrast to asset pricing, they do not aim to find interpretable factors that explain anomalies in excess asset returns. Our LLM-based approach offers an alternative interpretation through analysis reports.
Figure 2: Visualization of the keywords in the titles of news articles on the days over the highly complicated first two-year span in our dataset when the market return is positive (left) and negative (right).
Large Language Model Agents for Finance.
@@
LLM agents possess powerful emergent capabilities, such as reasoning, planning, and tool-using Achiam et al. (2023). The core of LLM agent programming lies in prompting, which employs contextual hinting text to regulate the output of LLM Liu et al. (2023). Several prompting strategies have been proposed. Chain-of-Thoughts (CoT) Wei et al. (2022) encourages the agent to reason in a step-by-step manner. Yao et al. (2022) introduced the ReAct prompt, enabling the agent to refine its output based on the results of previous attempts. It allows the agent to use external tools, such as databases and search engines. Memory is another crucial component of LLM agents. Hu et al. (2023) introduced databases as symbolic memories. Packer et al. (2023) stores dialogues in both long- and short-term memory, analogously to operating systems. Cheng & Chin (2024) developed an agent capable of making “investment” decisions on social science time series based on input news, reports, etc., and knowledge base, as well as the Internet. We focus on using the agent to simulate discretionary investment decision-making to synergize qualitative and quantitative asset pricing. LLM Agent has been applied for stock prediction, Koa et al. (2024); Ding et al. (2024); Zhang et al. (2024); Fatemi & Hu (2024) proposed multi-LLM agents to predict the movement direction of specific stocks while we focus on the fair pricing of all stocks in the market.
@@ 3 Method
Given a state vector $V_{\tau,a}$ at a time point $\tau\in\{0,1,2,...\}$, which represents the current status of the market, society, and an asset $a$, an asset pricing model predicts the excess returns $r_{\tau+1,a}$ of the asset at the subsequent time point, expressed as $P(r_{\tau+1,a}|V_{\tau,a})$. In our study, each time point corresponds to one day. In traditional factor-based methods, the state $V_{\tau,a}\in\mathcal{N_{F}}^{N}$ is a vector composed of $N_{F}$ factors that are manually derived from economic indicators, market data, asset characteristics, etc. For instance, the market excess return, the performance disparity between small and large firms, and the difference between high and low book-to-market companies in the Fama-French 3-factor model Fama & French (1992). Recently, Bybee et al. (2021) demonstrated that a collection of business news can serve as an alternative representation of macroeconomic conditions, while Bybee et al. (2023) employs LDA to extract news characteristics as economic predictors for pricing. Building on this idea, we use the average embedding of analysis reports that mine values from the news as a proxy for the society, economic, and market states.
Business news in major media outlets like the WSJ carries important market information, however, they typically restrict their interpretations and opinions, leaving room for discretionary analysis. It is crucial to understand that business events are often interrelated.
@@
Based on these observations, we introduce an LLM agent asset pricing model, utilizing an LLM agent with long-term memory of domain knowledge and historical news analysis to iteratively analyze the input news and generate the analysis report, as detailed in Section 3.1. Subsequently, we combine these qualitative analysis reports and quantitative manual factors to feed into our hybrid asset pricing network in Section 3.2.
Figure 3: The demonstration of our hybrid asset pricing network. The purple boxes mark the computational components. Yellow boxes are data, and the circled plus symbol means concatenation. The MSE loss is computed with predicted returns feedback to update the network.
@@ 3.1 Discretionary Analysis with LLM Agent
The agent utilizes the latest news $x_{t}$ at time $t$ (e.g., a WSJ article published at 9:32 AM on 6 June 2020), along with a note $n_{t}$ on macroeconomics and market trends, to generate an analysis report $R_{t}$. The note $n_{t}$ is initialized with a macroeconomics summary $n_{0}$ produced by an LLM, before a cut-off date $d_{k}$. It offers necessary macroscopic context on economic and societal trends not directly available from the news or the memory. The note is then iteratively updated to $n_{t^{\prime}}$ with the new analysis report $R_{t}$ to keep the context up-to-date, and we also prompt the agent to document investment ideas and market thoughts in the notes to provide a short-term background such as the trends on the market, long-term research opportunities to watch. To ensure the note is continuously updated without missing information, the dataset in our study starts from $d_{k}+1$, immediately following the cut-off date.
@@
The analysis process begins with generating a refined news item $x^{\prime}_{t}$ that summarizes key information from the raw input $x_{t}$. This step helps control the input length and standardizes the format and style. The refined news $x^{\prime}_{t}$ and the note $n_{t}$ are then combined to form an input $I_{t}$ for the agent. The agent will determine if the news contains investment information: if not, it will be skipped; otherwise, an initial analysis report $R_{t}^{0}$ will be created. The report undergoes iterative refinement over $N$ rounds. In each round $i$, the report $R_{t}^{i-1}$ is used to query an external memory $M^{t}$, a vector database initialized with the SocioDojo knowledge base Cheng & Chin (2024), which includes textbooks, encyclopedias, and academic journals in fields such as economics, finance, business, politics, and sociology. We use BGE Xiao et al. (2023) as the embedding model $f_{e}$, which maps text to a vector $e\in\mathcal{R}^{d_{emb}}$ for querying the memory. This choice is based on the MTEB leaderboard Muennighoff et al. (2022), where we selected the best retrieval model considering performance, model size, and embedding vector length. In each round $i$, the top-$K$ most relevant items $\{m^{t,i}_{j}\}_{j=1}^{K}\subset M^{t}$ are retrieved and provided to the agent along with the report $R^{i-1}_{t}$ to produce the refined report $R_{t}^{i}$. The report $R_{t}^{N}$ generated after the $N$-th round is used as the final analysis report $R_{t}$ for the news $x_{t}$ and to update the note as $n_{t^{\prime}}$. Then it is inserted into the memory $M^{t}$ for future reference and pricing, updating the memory to $M^{t^{\prime}}$.
@@
The pricing network will utilize the analysis reports $\{R_{t^{d}_{i}}\}_{i=1}^{N_{d}}$ of all filtered news $\{x_{t^{d}_{i}}\}_{i=1}^{N_{d}}$ for a given day $d$, where $N_{d}$ represents the number of filtered news items on day $d$. Figure 1 provides an overview of the entire analysis process. The prompts employed by our agent are detailed in Appendix E. In Section 4.4, we conduct experiments on our agent design and the impact of $N$ and $K$.
@@ 3.2 Hybrid Asset Pricing Network
where $\kappa(L,i)$ is an exponential decay kernel defined as $\frac{\eta^{L-i}}{\sum_{j=1}^{L}\eta^{L-j}}$. The decay coefficient is denoted as $10<\eta<$ 1. We form a raw hybrid state $h_{d,a}=[s_{d};v_{d,a}]$ by concatenating the smoothed daily state $s_{d}$ with a vector $v_{d,a}\in\mathcal{R}^{N_{F}}$ of $N_{F}$ manual-constructed financial economic factors. The asset $a$ is indexed by a permanent number (permno) from the Center for Research in Security Prices (CRSP)  [^1] https://www.crsp.org/ database. The hybrid state is subsequently downsampled by $h^{\prime}_{d,a}=\sigma(W_{S}h_{d,a})$, where $\sigma$ denotes the ReLU function and $W_{S}\in\mathcal{R}^{d_{model}\times(d_{emb}+N_{F})}$ is a parameter matrix.
To capture the asset-specific loading to the hybrid state, especially to the asset-agnostic $s_{d}$, we define an asset embedding $E\in\mathcal{R}^{N_{A}\times d_{model}}$, which can be looked up via the permnos of the assets. Here, $N_{A}$ denotes the total number of assets and $d_{model}$ is the dimension of the embedding. We then concatenate the asset embedding $E_{a}$ with the downsampled hybrid state to form $\hat{h}_{d,a}=[h_{d,a}^{\prime};\sigma(E_{a})]$, the asset-specfic hybrid state for $a$.
@@
$\begin{split}\arg\min_{\theta}\frac{1}{N_{A}N_{D}}\sum_{d,a}(r_{d+1,a}-\hat{r}_{d+1,a})^{2},\ where\ r_{d+1,a}=f_{H}(h_{d,a};\theta)\end{split}$
Where $N_{D}$ denotes the number of days in the training set. The model is trained for $T$ episodes with a batch size of $B$. We initially pre-train this hybrid asset pricing network $f_{H}$ to make use of the historical factor data available before the beginning of the news dataset. During this pre-training phase, a placeholder embedding (such as the embedding for the word ”Null”) is utilized.